# Both Text and Images Leaked! A Systematic Analysis of Multimodal LLM Data Contamination

### v1

@@ Abstract
The rapid progression of multimodal large language models (MLLMs) has demonstrated superior performance on various multimodal benchmarks. However, the issue of data contamination during training creates challenges in performance evaluation and comparison. While numerous methods exist for detecting dataset contamination in large language models (LLMs), they are less effective for MLLMs due to their various modalities and multiple training phases. In this study, we introduce a multimodal data contamination detection framework, MM-Detect, designed for MLLMs. Our experimental results indicate that MM-Detect is sensitive to varying degrees of contamination and can highlight significant performance improvements due to leakage of the training set of multimodal benchmarks. Furthermore, We also explore the possibility of contamination originating from the pre-training phase of LLMs used by MLLMs and the fine-tuning phase of MLLMs, offering new insights into the stages at which contamination may be introduced.
@@ 1 Introduction
While numerous works in the field of LLMs have proposed methods for detecting data contamination (Yeom et al., 2018; Deng et al., 2024; Dong et al., 2024), MLLMs, due to their various modalities and multiple training phases (Liu et al., 2023a; Li et al., 2023), face limitations when applying these methods. Therefore, there is a pressing need for a more suitable multimodal contamination detection framework specifically tailored for MLLMs.
In this study, we carry out a systematic analysis of multimodal data contamination. Initially, we define Multimodal Data Contamination, as it pertains to the modality of data sources exposed to the MLLMs, into two categories: Unimodal Contamination and Cross-modal Contamination. Subsequently, we unveil a detection framework for multimodal data contamination, MM-Detect, which incorporates two methods, Option Order Sensitivity Test and Slot Guessing for Perturbation Caption, designed to handle two common types of Visual Question Answering (VQA) tasks: multiple-choice and caption-based questions, respectively. Then, applying MM-Detect on eleven widely-used MLLMs across five prevalent VQA datasets, we observe that both open-source and proprietary MLLMs do exhibit contamination, with the degree of contamination varying across different models.
To corroborate the validity and sensitivity of our approach, we deliberately induce contamination in MLLMs, thus simulating feasible real-world scenarios. Experimental results indicate that our method proves to be quite effective and sensitive in identifying varying degrees of contamination. Interestingly, our findings reveal that not only does leakage in the multimodal benchmark test set play a role, but the training set can also contribute significantly to enhancing the model’s performance.
@@
  • We formulate the definition for multimodal contamination detection and present the MM-Detect framework, comprising two innovative methods specifically designed for effective contamination detection in MLLMs.
  • We demonstrate that leakage from multimodal benchmark data can significantly boost the model’s performance on test sets, with this enhancement intensifying as the degree of contamination increases.
@@ 2 Preliminaries
Figure 1: A description of Multimodal Data Contamination (left) and the overview of proposed MM-Detect framework (right).
We formally define the problem of contamination in multimodal benchmark datasets and describe the unique challenges in multimodal contamination detection.
@@ 2.1 Definition of Multimodal Data Contamination
In contrast to single-modal contamination, multimodal contamination may arise from both unimodal and multimodal data sources, as depicted in Figure 1. The training data for MLLMs generally consists of pure text pre-training data $D_{\text{pretrain}}$ and multimodal alignment or instruction-following data $D_{\text{vision}}$. Consider an instance $(x,i,y)$ from a benchmark dataset $D$, where $x$ represents the text input, $i$ is the image input, and $y$ is the label. Data contamination in multimodal models can be categorized into the following two cases:
@@ 2.2 Challenges in Multimodal Detection
Despite the prevalence of unimodal detection methods, their application in multimodal scenarios often encounters difficulties. For example, retrieval-based methods Brown et al. (2020); Touvron et al. (2023a) attempt to detect contamination by retrieving large-scale corpora used for model training. Yet, they struggle when retrieving multimodal information. Similarly, logits-based methods (Shi et al., 2024; Yeom et al., 2018) rely on observing the distribution of low-probability tokens in model outputs, but the disparity in token probability distributions is less pronounced in instruction-tuned models. Masking-based methods Deng et al. (2024), which assess training contamination by evaluating a model’s ability to predict specific missing or masked text, face challenges when images in multimodal samples provide clues, leading to overestimated contamination detection. Finally, comparison-based methods Dong et al. (2024) that measure contamination by comparing model outputs with benchmark data prove to be ineffective for image caption tasks due to low output similarity. To validate these inefficiencies, we conducted experiments with compelling results, which are detailed in Appendix A.
@@
Another challenge in detecting contamination in multimodal models is the multi-stage nature of their training (Yin et al., 2023). Each stage may be subject to data contamination. Initially, the pretraining corpus could contain the textual components of questions from benchmark test samples. Moreover, in certain native multimodal model training (Reid et al., 2024), test samples may be entirely exposed. Subsequently, during multimodal fine-tuning, the model may utilize training samples of some benchmarks, leading to skewed performance improvements. Furthermore, some models employ extensive mixed image-text data from the internet for modality alignment training (Lin et al., 2023; Bai et al., 2023b), potentially introducing additional contamination. Given these challenges, the development of an effective detection framework for multimodal contamination becomes an urgent need.
@@ 3 MM-Detect
  • The first step is to generate perturbed datasets using two innovative methods: the Option Order Sensitivity Test (Section 3.1) and the Slot Guessing for Perturbation Captions (Section 3.2), tailored for evaluating multiple-choice and image captioning tasks, respectively.
  • The second step involves the application of predefined metrics to detect contamination (Section 3.3), conducting thorough analyses at both the dataset and instance levels.
@@ 3.1 Option Order Sensitivity Test
Let $D$ represent a dataset comprising $n$ datapoints. For each datapoint $d_{i}$, where $i\in\{1,\ldots,n\}$, there is a question symbolized by $Q_{i}$, an image represented by $I_{i}$, and a list of choices is denoted by $A_{i}$, such that $A_{i}=\{a_{i}^{1},a_{i}^{2},\ldots,a_{i}^{m}\}$ and $m$ is the number of choices for that datapoint. The correct answer is symbolized as $a_{i}^{c}$, where $a_{i}^{c}\in A$. The list $A$ is randomly shuffled to generate $A^{\prime}$, ensuring that the index of the correct answer $a_{i}^{c}$ in $A^{\prime}$ differs from its index in $A$, thereby altering the correct answer’s position. The final prompts, both before and after the shuffling, are the concatenation of the image, question, and choices:
$P=\text{Concat}(I,Q,A),P^{\prime}=\text{Concat}(I,Q,A^{\prime}),$
where $P$ and $P^{\prime}$ are the prompts fed into the model, and $Q$ and $I$ remain constant.
@@ 3.2 Slot Guessing for Perturbation Caption
Figure 3: An example of Slot Guessing for Perturbation Caption.
This method is based on the intuition that if a model can predict a missing part of a sentence but fails with the back-translated version (from English to Chinese, then back to English), it likely indicates that the model has encountered the original sentence during training.
@@
Let $D$ be a dataset containing $n$ datapoints. For each datapoint $d_{i}$, $i\in\{1,\ldots,n\}$, there is a corresponding caption $S_{i}$ describing the image features. We first apply a back-translation function [^1] We use Google-Translate API for Python to implement the back-translation. to $S_{i}$:
$S^{\prime}=f_{\text{back-translate}}(S)$
to obtain the back-translated sentence $S^{\prime}$. Next, we perform keyword extraction [^2] We employ the Stanford POS Tagger (Toutanvoa & Manning, 2000), targeting nouns, adjectives, or verbs, as they encapsulate the sentences’ core meaning. on both $S$ and $S^{\prime}$:
$K=f_{\text{keyword}}(S),K^{\prime}=f_{\text{keyword}}(S^{\prime}),$
where $K$ and $K^{\prime}$ are the keywords extracted from $S$ and $S^{\prime}$, respectively. We then use a masking function $f_{\text{mask}}$ to replace the keywords in the sentences with [MASK]:
$S_{\text{mask}}=f_{\text{mask}}(S,K),S^{\prime}_{\text{mask}}=f_{\text{mask}}(S^{\prime},K^{\prime}).$
The final prompt can be represented as the concatenation of the image, the instruction and the masked sentence:
$P=\text{Concat}(I,Q,S_{\text{mask}}),P^{\prime}=\text{Concat}(I,Q,S^{\prime}_{\text{mask}}),$
where $I$ is the image and $Q$ is the instruction guiding the model to complete the mask word prediction task.
@@ 3.3 Detection Metrics
Having introduced two detection methods, we now delineate the metrics for the detection pipeline, which consists of two primary steps.
Step 1: Benchmark Atomic Metrics Calculation.
This step assesses the model’s performance on benchmark $D$ before and after perturbation. We denote the correct rate (CR) and perturbed correct rate (PCR) uniformly for both Option Order Sensitivity Test (using Accuracy) and Back-Translation Keyword Guessing (using Exact Match). Here, $N$ and $N^{\prime}$ are the counts of correct answers before and after perturbation, respectively. They are calculated as:
@@
Dataset Level. We evaluate the reduction in atomic metrics, denoted as $\Delta$:
@@
Instance Level. Despite a non-significant or positive $\Delta$, contamination may still occur at the instance level, as some instances may still have been unintentionally included during training. To identify such instances, we compute $X$, the count of cases where the model provided correct answers before perturbation but incorrect answers after. The instance leakage metric $IL$ is then obtained by dividing $X$ by the dataset size:
$\text{IL}=\frac{X}{|D|},$
where a larger $IL$ indicates a higher likelihood of instance leakage. For further details on our atomic metrics’ computation, please refer to Appendix B.

### v2

# Both Text and Images Leaked! A Systematic Analysis of Data Contamination in Multimodal LLM

@@ Abstract
The rapid progression of multimodal large language models (MLLMs) has demonstrated superior performance on various multimodal benchmarks. However, the issue of data contamination during training creates challenges in performance evaluation and comparison. While numerous methods exist for detecting models’ contamination in large language models (LLMs), they are less effective for MLLMs due to their various modalities and multiple training phases. In this study, we introduce a multimodal data contamination detection framework, MM-Detect, designed for MLLMs. Our experimental results indicate that MM-Detect is quite effective and sensitive in identifying varying degrees of contamination, and can highlight significant performance improvements due to the leakage of multimodal benchmark training sets. Furthermore, we explore whether the contamination originates from the base LLMs used by MLLMs or the multimodal training phase, providing new insights into the stages at which contamination may be introduced.
{CJK}
UTF8gbsn
Both Text and Images Leaked! A Systematic Analysis of Data Contamination in Multimodal LLM
[…]
@@ 1 Introduction
While numerous works in the field of LLMs have proposed methods for detecting data contamination (Yeom et al., 2018; Deng et al., 2024; Dong et al., 2024), MLLMs, due to their various modalities and multiple training phases (Liu et al., 2023a; Li et al., 2023), face limitations when applying these methods. Therefore, there is a pressing need for a multimodal contamination detection framework specifically tailored for MLLMs.
Figure 1: A description of Multimodal Data Contamination, which can originate from various training stages.
In this study, we carry out a systematic analysis of multimodal data contamination. Initially, we define Multimodal Data Contamination, as it pertains to the modality of data sources exposed to the MLLMs, into two categories: Unimodal Contamination and Cross-modal Contamination, as illustrated in Figure 1. Subsequently, we unveil a detection framework, MM-Detect, which incorporates two methods, Option Order Sensitivity Test and Slot Guessing for Perturbation Caption, designed to handle two common types of Visual Question Answering (VQA) tasks: multiple-choice and caption-based questions, respectively.
To corroborate the validity and sensitivity of our approach, we deliberately induce contamination in MLLMs, thus simulating feasible real-world scenarios. Experimental results indicate that our method proves to be quite effective and sensitive in identifying varying degrees of contamination. Interestingly, our findings reveal that not only does leakage in the multimodal benchmark test set play a role, but the training set can also contribute significantly to enhancing the model’s performance. Then, applying MM-Detect on twelve widely-used MLLMs across five prevalent VQA datasets, we find that both open-source and proprietary MLLMs exhibit contamination with varying degrees.
@@
  • We formulate the definition for multimodal contamination and present the MM-Detect framework, comprising two innovative methods specifically designed for effective contamination detection in MLLMs.
  • We demonstrate that leakage from multimodal benchmark data—whether from the training, validation, or test sets—can significantly enhance models’ performance, with this performance gain intensifying as the degree of contamination increases.
@@ 2 Preliminaries
We formally define the problem of multimodal data contamination and outline the unique challenges associated with its detection.
@@ 2.1 Definition of Multimodal Data Contamination
In contrast to single-modal contamination, multimodal contamination may arise from both unimodal and multimodal data sources, as depicted in Figure 1. The training data for MLLMs generally consists of pure text pre-training data $D_{\text{pretrain}}$ and multimodal alignment or instruction-following data $D_{\text{vision}}$. Consider an instance $(x,i,y)$ from a benchmark dataset $D$, where $x$ represents the text input, $i$ is the image input, and $y$ is the label. Data contamination in MLLMs can be categorized into the following two cases:
@@ 2.2 Challenges in Multimodal Detection
Despite the prevalence of unimodal detection methods, their application in multimodal scenarios often encounters difficulties. For example, retrieval-based methods Brown et al. (2020); Touvron et al. (2023a) attempt to detect contamination by retrieving large-scale corpora used for model training. Yet, they struggle when retrieving multimodal information. Similarly, logits-based methods (Shi et al., 2024; Yeom et al., 2018) rely on observing the distribution of low-probability tokens in model outputs, but the disparity in token probability distributions is less pronounced in instruction-tuned MLLMs. Masking-based methods Deng et al. (2024), which assess training contamination by evaluating a model’s ability to predict specific missing or masked text, face challenges when images in multimodal samples provide clues, leading to overestimated contamination detection. Finally, comparison-based methods Dong et al. (2024) that measure contamination by comparing model outputs with benchmark data prove to be ineffective for image caption tasks due to low output similarity. To validate these inefficiencies, we have conducted comprehensive experiments with compelling results, which are detailed in Appendix A.
Figure 2: The overview of proposed MM-Detect framework.
@@
Another challenge in detecting contamination in MLLMs is the multi-stage nature of their training (Yin et al., 2023). Each stage may be subject to data contamination. 1) Initially, the pretraining corpus could contain the textual components of questions from benchmark samples. Moreover, in certain native multimodal model training (Reid et al., 2024), samples may be entirely exposed. 2) Subsequently, during multimodal fine-tuning, the model may utilize training samples of some benchmarks, leading to skewed performance improvements. 3) Furthermore, some models employ extensive mixed image-text data from the internet for modality alignment training (Lin et al., 2023; Bai et al., 2023b), potentially introducing additional contamination. Given the challenges, the development of an effective detection framework for multimodal contamination becomes an urgent need.
@@ 3 MM-Detect
  • The first step is to generate perturbed datasets using two innovative methods: the Option Order Sensitivity Test3.1) and the Slot Guessing for Perturbation Captions3.2), tailored for multiple-choice and image captioning tasks, respectively.
  • The second step involves the application of predefined metrics to detect contamination (§3.3), conducting thorough analyses at both the dataset and instance levels.
@@ 3.1 Option Order Sensitivity Test
Let $D$ be a dataset consisting of $n$ datapoints. Each datapoint $d_{i}$ ($i\in\{1,\ldots,n\}$) comprises a question $Q_{i}$, an associated image $I_{i}$, and a set of answer choices $A_{i}=\{a_{i}^{1},a_{i}^{2},\ldots,a_{i}^{m}\}$, where $m$ is the number of choices and the correct answer is denoted by $a_{i}^{c}$.
To introduce positional variation, the set $A_{i}$ is randomly shuffled to obtain a new set $A_{i}^{\prime}$, ensuring that the index of the correct answer $a_{i}^{c}$ in $A_{i}^{\prime}$ differs from its original position in $A_{i}$. The final prompts, before and after shuffling, are constructed by concatenating the image, question and choices:
$P=\text{Concat}(I_{i},Q_{i},A_{i}),$
$P^{\prime}=\text{Concat}(I_{i},Q_{i},A_{i}^{\prime}),$
where $P$ and $P^{\prime}$ are the inputs to the model, and $Q_{i}$ and $I_{i}$ remain unchanged throughout this process.
@@ 3.2 Slot Guessing for Perturbation Caption
This method is based on the intuition that if a model can predict a missing and important part of a sentence but fails with the back-translated version (from English to Chinese, then back to English), it likely indicates that the model has encountered the original sentence during training.
Figure 4: A simple example shows the procedure.
@@
Let $D$ be a dataset containing $n$ datapoints. Each datapoint $d_{i}$ ($i\in\{1,\ldots,n\}$) consists of an image-caption pair, where the caption $S_{i}$ describes the visual features of the corresponding image $I_{i}$. We first apply a back-translation function [^1] We use the Google Translate API for Python to implement back-translation. to $S_{i}$:
$S^{\prime}_{i}=f_{\text{back-translate}}(S_{i}).$
resulting in a paraphrased version $S^{\prime}_{i}$. Next, we perform keyword extraction [^2] We employ the Stanford POS Tagger (Toutanvoa and Manning, 2000), targeting nouns, adjectives, and verbs, as they encapsulate the core meaning of the sentences. on both $S_{i}$ and $S^{\prime}_{i}$:
$K_{i}=f_{\text{keyword}}(S_{i}),\quad K^{\prime}_{i}=f_{\text{keyword}}(S^{\prime}_{i}),$
where $K_{i}$ and $K^{\prime}_{i}$ denote the extracted keywords from $S_{i}$ and $S^{\prime}_{i}$, respectively. We then apply a masking function $f_{\text{mask}}$ to replace the extracted keywords with a placeholder token [MASK]:
$S_{i,\text{mask}}=f_{\text{mask}}(S_{i},K_{i}),~{}sS^{\prime}_{i,\text{mask}}=f_{\text{mask}}(S^{\prime}_{i},K^{\prime}_{i}).$
The final prompt guiding the model to complete the masked-word prediction can be represented as:
$P_{i}=\text{Concat}(I_{i},Q_{i},S_{i,\text{mask}}),$
$P^{\prime}_{i}=\text{Concat}(I_{i},Q_{i},S^{\prime}_{i,\text{mask}}).$
@@ 3.3 Detection Metrics
Having introduced two detection methods, we now delineate the atomic metrics for the detection pipeline, which consists of two primary steps.
Step 1: Correct Rate Calculation.
This step assesses the model’s performance on benchmark $D$ before and after perturbation. We denote the correct rate (CR) and perturbed correct rate (PCR) uniformly for both Option Order Sensitivity Test (using Accuracy) and Slot Guessing (using Exact Match). Here, $N$ and $N^{\prime}$ are the counts of correct answers before and after perturbation, respectively. They are calculated as:
@@
Dataset Level Metric. We evaluate the reduction in atomic metrics, denoted as $\Delta$:
@@
Instance Level Metric. Despite a non-significant or positive $\Delta$, contamination may still occur at the instance level, as some instances may still have been unintentionally included during training. To identify such instances, we compute $X$, the count of cases where the model provided correct answers before perturbation but incorrect answers after. The instance leakage metric $\Phi$ is then obtained by dividing $X$ by the dataset size:
$\Phi=\frac{X}{|D|},$
where a larger $\Phi$ indicates a higher likelihood of instance leakage.
Contamination Degree.
We define the contamination degree solely based on $\Delta$, capturing how much performance changes from $D$ to $D_{\text{pert}}$. The threshold for each severity level (Minor, Partial, Severe) is derived from experimental findings in §4. The entire pipeline, including how these levels are assigned, is summarized in Algorithm 1.
Algorithm 1 MM-Detect Framework
0:  Benchmark dataset $D$, Model $M$
1:  Define contamination degree $\mathcal{C}_{\text{Minor}}$, $\mathcal{C}_{\text{Partial}}$, $\mathcal{C}_{\text{Severe}}$
2:  if $D$ is multiple-choice then
3:     Generate perturbed set $D_{\text{pert}}$ via §3.1
4:  else
5:     Generate perturbed set $D_{\text{pert}}$ via §3.2
6:  end if
7:  Compute $CR$, $PCR$, $\Delta$, $\Phi$ using §3.3
8:  if multiple-choice then
9:     $\mathcal{C}\leftarrow\begin{cases}\mathcal{C}_{\text{Minor}},&\Delta\in(-1.6,-0.2]\\ \mathcal{C}_{\text{Partial}},&\Delta\in(-2.9,-1.6]\\ \mathcal{C}_{\text{Severe}},&\Delta\leq-2.9\end{cases}$
10:  else
11:     $\mathcal{C}\leftarrow\begin{cases}\mathcal{C}_{\text{Minor}},&\Delta\in(-2.4,-1.1]\\ \mathcal{C}_{\text{Partial}},&\Delta\in(-5.0,-2.4]\\ \mathcal{C}_{\text{Severe}},&\Delta\leq-5.0\end{cases}$
12:  end if
12:  $CR$, $PCR$, $\Delta$, $\Phi$, $\mathcal{C}$
4 Intentional Contamination
In this section, we address three research questions through intentional contamination:
RQ1: Is MM-Detect an effective detector?
RQ2: How sensitive is MM-Detect?
RQ3: Does training set leakage cause bias?
To answer these questions, we adopt the LLaVA framework and train several 7B-parameter models using intentionally contaminated downstream task data during the visual instruction tuning phase. We then evaluate the degree of contamination in each model. The contamination data are identical to those described in §5.1.
4.1 MM-Detect is An Effective Detector
We reproduced the LLaVA-1.5-7B experiment to obtain a baseline model without contamination. Recognizing that contamination can occur anywhere in the training data, we inserted contaminated samples into the visual instruction tuning dataset ($D_{\text{tuning}}$) at three positions, early, mid, and late, creating two groups of contaminated training sets using 1340 ScienceQA test samples or 1000 NoCaps validation samples. Corresponding models, termed Early Cont., Mid Cont., and Late Cont., were then trained for comparison with the baseline.
Models ScienceQA Test Set NoCaps Val. Set
CR PCR $\Delta$ CR PCR $\Delta$
Baseline 61.4 61.5 0.01 33.0 32.1 -0.9
Early Cont. 71.5 68.1 -3.4 37.5 32.0 -5.5
Mid Cont. 69.4 67.3 -2.1 38.5 35.1 -3.4
Late Cont. 70.2 66.9 -3.3 38.7 32.6 -6.1
Table 1: Detection results on contamination using the ScienceQA test set and NoCaps validation set.
Table 1 shows that incorporating contaminated data during training increases both the model’s performance and its sensitivity to perturbations. Compared with the baseline, ScienceQA-contaminated models exhibit average increases in CR and PCR of 9.0% and 5.9%, respectively, while NoCaps-contaminated models show increases of 5.2% and 1.1%. Moreover, all contaminated models demonstrate a marked decrease in $\Delta$, confirming that MM-Detect effectively identifies data contamination. We define the average $\Delta$ across these models as the threshold for severe contamination ($\mathcal{C}_{\text{Severe}}$); specifically, $\Delta\leq-2.9$ indicates severe leakage in the multiple-choice dataset, and $\Delta\leq-5$ indicates severe leakage in the caption dataset.
4.2 MM-Detect is Sensitive and Fine-grained
We evaluated MM-Detect’s sensitivity by varying leakage levels in the training set. Using the fully contaminated model as our baseline, we trained additional models with moderate and minimal contamination, by inserting reduced amounts (10% and 50%) of contaminated data at the late position of the training set, to assess leakage impact.
Figure 5: Performance and atomic metrics evaluated under varying leakage levels on the ScienceQA test set and NoCaps validation set.
As illustrated in Figure 5, increasing contamination from 10% to 50% to 100% results in corresponding increases in CR and PCR, alongside progressively larger $\Delta$ values. We define the $\Delta$ for 50% contamination as the threshold for partial leakage ($\mathcal{C}_{\text{Partial}}$) and that for 10% contamination as the threshold for minor leakage ($\mathcal{C}_{\text{Minor}}$), as detailed in Algorithm 1. These findings confirm that our framework can accurately differentiate between varying leakage levels in benchmark datasets.
4.3 Training Set Leakage Leads to Unfairness
We investigated whether training set leakage biases evaluations by comparing models trained with and without benchmark data contamination. For the ScienceQA experiment, we appended 2000 ScienceQA training samples to the training dataset, creating a contaminated model. For the COCO experiment, we removed the COCO-Caption2017 training data from the original training dataset, resulting in a model without leakage.
Model Dataset CR PCR $\Delta$
Clean ScienceQA 61.4 61.5 0.01
Leaked ScienceQA 64.3 63.8 -0.5
Clean COCO-Caption2017 32.5 31.9 -0.6
Leaked COCO-Caption2017 38.1 34.9 -3.2
Table 2: Performance of models trained without (Clean) and with (Leaked) training set contamination.
Table 2 compares the models’ performance. On the ScienceQA test set, the contaminated model outperforms the clean model by 2.9% in CR and 2.3% in PCR, with a $\Delta$ of -0.5, meeting the minor leakage threshold ($\mathcal{C}_{\text{Minor}}$). On the COCO-Caption2017 validation set, the model with COCO data shows a $\Delta$ of -3.2, exceeding the partial leakage threshold ($\mathcal{C}_{\text{Partial}}$). The results indicate that training set leakage inflates performance and that MM-Detect can effectively detect it likewise.
Takeaways Both training and test set leakage can result in unfairness, and the degree of contamination can be detected through MM-Detect effectively.